OpenStudy (anonymous):

The human population of a small pacific island satisfies the logistic law dx/dt = kx - λx^2, with k=.04, λ=2(10^-7), and time t measured in years. Find a formula for the population in future years.

OpenStudy (anonymous):

I got this far: please help!

OpenStudy (anonymous):

also I have the answer. It should be:

OpenStudy (dumbcow):

what is initial population?

OpenStudy (anonymous):

50,000

OpenStudy (dumbcow):

ok i think your "B" coefficient is wrong I think its easier or at least less confusing if you solve the diff equ with the constants then plug in the values later I get A = 1/k B = L/k

OpenStudy (dumbcow):

\[\int\limits \frac{dx}{x(k-\lambda x)} = \int\limits dt\] \[\int\limits \frac{A}{x} + \frac{B}{k - \lambda x} = \int\limits dt\] \[Ak + (B - \lambda A)x = 1\] \[A = \frac{1}{k}\] \[B = \lambda A = \frac{\lambda}{k}\]

OpenStudy (dumbcow):

after integrating you should get \[\frac{1}{k}[\ln x - \ln (k -\lambda x)] = t + C\] combine logs and raise everything as power of e \[\frac{x}{k - \lambda x} = C e^{kt}\] I would solve for constant of integration here using initial value ,50,000 or 5(10^4) \[C = \frac{5(10^4)}{4(10^{-2}) - 2(10^{-7})*5(10^4)} = \frac{5 (10^6)}{3}\] Using algebra and solving for x \[x = \frac{C k e^{kt}}{1 + C \lambda e^{kt}}\] plug in values , k, lambda, C ... simplify

OpenStudy (anonymous):

thanks so much but can you explain why the book answer has 3 + in the denominator

OpenStudy (anonymous):

nevermind I got it THANK YOU